OpenStudy (anonymous):

rewrite the following in standard form? x^2+14x+44y+313=0 3y^2+6y+108x-969=0 x^2-12x-48y+372=0

OpenStudy (anonymous):

Do you know what is the standard form of a parabola?

OpenStudy (anonymous):

No, I'm really lost :(

OpenStudy (anonymous):

The standard form is y=ax^2 +bx +c

OpenStudy (anonymous):

Where a,b and c are real numbers. (any number)

OpenStudy (anonymous):

Ok, so could you just show me how to one of them first as an example?

OpenStudy (anonymous):

x^2+14x+44y+313=0 So I will help you with the above. 44y has to be taken to the other side of the equation. Thus we subtract 44y from both sides. x^2+14x+31=-44y

OpenStudy (anonymous):

Have to go for a while now

OpenStudy (anonymous):

Oh, so thats it?

OpenStudy (anonymous):

x^2+14x+31=-44y Just dived it by -44 to get the standard form. It will look ugly :)

OpenStudy (anonymous):

The goal is that the equation should look like y=something something something